OpenStudy (mayankdevnani):

Another limit question. Options :- A) 9 p (ln 4) B) 3 p (ln 4)^3 C) 12 p (ln 4)^3 D) 27 p (ln 4)^2

OpenStudy (mayankdevnani):

\[\huge \bf \lim_{x \rightarrow 0} \frac{(4^x-1)^3}{\sin(\frac{x}{p})\ln(1+\frac{x^2}{3})}\]

OpenStudy (freckles):

that one doesn't look easy to please but when we plug in 0 we get 0/0 so my first thoughts are to do the l'hospital thing and that will look ugly though I will have to do a bit more thinking on this one

OpenStudy (perl):

I would use Lhopitals

OpenStudy (mayankdevnani):

can you please show your next step using L HOPITAL's rule ?

OpenStudy (mayankdevnani):

look at options ! @freckles and @perl

OpenStudy (perl):

$$ \LARGE \rm{ \lim_{x \rightarrow 0}~ \frac{(4^x-1)^3}{\sin(\frac{x}{p})\ln(1+\frac{x^2}{3})} \\~\\\ \Rightarrow ^{Lhopitals} \\ \lim_{x \rightarrow 0}~\frac{ 6(4^x-1)^24^x~ln(2)} { \frac{cos(\frac x p ) ln(1+\frac13 x^2)}{p}+\frac 23\frac{ sin(\frac xp )x}{(1+\frac 13 x^2)} } } $$

OpenStudy (mayankdevnani):

oh !! so lengthy ! don't think it will match the options !

OpenStudy (mayankdevnani):

i don't think this question will require L HOSTPITAL'S rule ? @perl and @freckles

OpenStudy (mayankdevnani):

any other simplest method ???

OpenStudy (perl):

you can try to take the logarithm of the limit

OpenStudy (perl):

suppose that limit is equal to L. now find ln(L)

OpenStudy (freckles):

well you can cheat I think I call this cheating anyways set p=1 so your options are now A) 9 (ln 4) B) 3 (ln 4)^3 C) 12 (ln 4)^3 D) 27 (ln 4)^2 where your problem is now \[\huge \bf \lim_{x \rightarrow 0} \frac{(4^x-1)^3}{\sin(x)\ln(1+\frac{x^2}{3})}\] and plug in a number like .001 and see which of the choices are closer to that output I don't like numerical methods for finding or approximating limits but that is all I got right now still trying to search for other options

OpenStudy (perl):

you can also brute force the options. plug in a value for p, and let x get very close to zero , say .0001. which of the choices match

OpenStudy (mayankdevnani):

i don't know how to solve logarithm limit ? I AM NEW LEARNER

OpenStudy (perl):

well i thought wolfram would be able to handle it, it can't :(

OpenStudy (mayankdevnani):

@freckles you will surprise that in my grade maximum students will solve this question ! without putting p=1 and all that

OpenStudy (freckles):

it can handle it without the p

OpenStudy (mayankdevnani):

how ?

OpenStudy (mayankdevnani):

`i am in hurry` please type fast

OpenStudy (freckles):

I was talking about wolfram

OpenStudy (mayankdevnani):

oh !

OpenStudy (freckles):

wait perl it can do it with the p

OpenStudy (mayankdevnani):

then i have to consult this question by teacher or my friends ! and then i will definitely post the solution !

OpenStudy (freckles):

lol I thought it was goes to 0 not infinity

OpenStudy (mayankdevnani):

BBYYEE.,.. @freckles @perl I HAVE TO GO NOW !!!

OpenStudy (perl):

i get it is 8

OpenStudy (perl):

with p = 1 , the limit is 8

OpenStudy (freckles):

@perl it x->0

OpenStudy (perl):

nevermind

OpenStudy (perl):

3 ln (4)^3 is the answer

OpenStudy (freckles):

hey @perl did you get l'hospital to work?

OpenStudy (freckles):

like did you actually complete that way?

OpenStudy (mayankdevnani):

hey @perl you are correct

OpenStudy (mayankdevnani):

can you please explain ????

OpenStudy (perl):

Applying Lhopitals once gave that more complicated expression. I used the calculator , let .0001 -> A , store it

OpenStudy (perl):

i put in x -> infinity , lol

OpenStudy (perl):

in wolfram, that was my mistake

OpenStudy (perl):

hmm, they say step by step solution, i wonder

OpenStudy (perl):

opens wolfram mathematica 10

OpenStudy (freckles):

hey i think I have something interesting

OpenStudy (freckles):

\[\lim_{x \rightarrow 0}\frac{(\frac{4^x-1}{x})^3}{\frac{\sin(x)}{x} \cdot \frac{\ln(1+\frac{x^2}{3})}{x^2}}\]

OpenStudy (freckles):

I divided top and bottom by x^3

OpenStudy (freckles):

top goes to [ln(4)]^3

OpenStudy (freckles):

sin(x)/x->1

OpenStudy (freckles):

and that last thingy should go to (1/3)

OpenStudy (freckles):

\[\lim_{x \rightarrow 0}\frac{(\frac{4^x-1}{x})^3}{\frac{\sin(x)}{x} \cdot \frac{\ln(1+\frac{x^2}{3})}{x^2}}=\frac{(\ln(4))^3}{1 \cdot \frac{1}{3}}\]

OpenStudy (mayankdevnani):